OpenStudy (anonymous):

One cookie recipe calls for 1/2 cup of flour, and a second recipe calls for 3/5 cup of flour. How much flour will a baker need to make both recipes? Answer 4/7 c. 1 1/10 c. 2/3 c. 1 1/5 c.

OpenStudy (anonymous):

@matricked

OpenStudy (anonymous):

Add the two together 1/2+3/5=11/10 cups

OpenStudy (anonymous):

3/5 +1/2 = .6+.5 =1.10=1 1/10 c.